# Pairs of Clark Unitary Operators on the Bidisk and their Taylor Joint Spectra

### v1

@@ Abstract.
We develop a Clark theory for commuting compressed shift operators on model spaces $K_{\phi}$ associated with inner functions $\phi$ on the bidisk, which exhibits both similarities and marked differences compared to the classical one-variable version. We first identify the adjoint of the embedding operator $J_{\alpha}\colon K_{\phi}\to L^{2}(\sigma_{\alpha})$ as a weighted Cauchy transform of the Clark measure $\sigma_{\alpha}$. Under natural assumptions, which generically include the case when $\phi$ is rational inner, we then use $J_{\alpha}^{*}$ to obtain commuting unitaries on $K_{\phi}$ which are often infinite-dimensional perturbations of compressed shift operators. Finally, we show that the Taylor joint spectrum of these Clark unitaries coincides with $\mathrm{supp}(\sigma_{\alpha})\subset\mathbb{T}^{2}$ when $\phi$ is a rational inner function.
@@ 1.1. Introduction
is a fundamental example of an isometry. It is defined on holomorphic functions $f=\sum_{k=0}^{\infty}a_{k}z^{k}$ in the disk satisfying $\|f\|^{2}_{H^{2}(\mathbb{D})}=\sum_{k}|a_{k}|^{2}<\infty$ as $T\colon f\mapsto z\cdot f$. The shift operator has been studied extensively over the last century, and many important theorems regarding the Hardy shift have become classical results in operator theory (see, for instance, [22, Chapters 3 and 4]). An important example is a well-known theorem of Beurling which characterizes the invariant subspaces of the shift. Namely, $\mathcal{M}\subset H^{2}(\mathbb{D})$ is a non-trivial closed invariant subspace for $T$ if and only if $\mathcal{M}=\phi H^{2}(\mathbb{D})$ for some inner function $\phi\colon\mathbb{D}\to\mathbb{C}$. Recall that $\phi$ is inner if $\phi\in H^{\infty}(\mathbb{D})$, the Banach algebra of bounded holomorphic functions in the disk, and its non-tangential limits satisfy $|\phi^{*}(\zeta)|=1$ for almost every $\zeta$ on the unit circle
@@
$S_{\phi}=P_{K_{\phi}}T,$
where $P_{K_{\phi}}\colon H^{2}(\mathbb{D})\to K_{\phi}$ denotes orthogonal projection. The ‘model’ terminology arises from Szőkefalvi-Nagy–Foiaş theory, which realizes a wide class of Hilbert space contractions as compressed shifts on model spaces [22, Chapter 9]. This fact provides a powerful motivation for studying compressions of the shift acting on model spaces, and also for investigating how properties of the underlying inner function $\phi$ are reflected in its operator theory. An important contribution in this direction was made by D.N. Clark, who observed [17] that the rank one perturbations of $S_{\phi}$ to a unitary operator form a one-parameter family $\{U_{\alpha}\}$ indexed by $\alpha\in\mathbb{T}$. He further identified the spectral measures of these unitaries as the positive singular Borel measures $\{\sigma_{\alpha}\}$ that satisfy
$\frac{1-|\phi(z)|^{2}}{|\alpha-\phi(z)|^{2}}=\int_{\mathbb{T}}P_{z}d\sigma_{\alpha}$ (1.1)
where $P_{z}$ denotes the Poisson kernel of the disk. These measures are often referred to as Clark measures , and their properties are intimately connected, through a unitary map $J_{\alpha}\colon K_{\phi}\to L^{2}(\sigma_{\alpha})$, with both the function theory of $\phi$ and the spectral theory of $U_{\alpha}$. For instance, $\sigma_{\alpha}$ is the spectral measure of $U_{\alpha}$, and a point in $\mathrm{supp}\,\sigma_{\alpha}\subset\mathbb{T}$ is an eigenvalue of $U_{\alpha}$ precisely if $\phi$ has a finite angular derivative at the corresponding point of the unit circle. We refer the reader to [16, 22] for accessible introductions to classical Clark theory, and to [32, 33, 21, 28] for a more varied accounting of recent progress (eg. including the case of non-inner $\phi$).
@@
and its distinguished boundary $\mathbb{T}^{2}=\{(\zeta_{1},\zeta_{2})\in\mathbb{C}^{2}\colon|\zeta_{j}|=1,\,j=1,2\}$. We say that a holomorphic function $f\colon\mathbb{D}^{2}\to\mathbb{C}$ belongs to $H^{2}(\mathbb{D}^{2})$ if $f=\sum_{k=0}^{\infty}\sum_{l=0}^{\infty}a_{k,l}z_{1}^{k}z_{2}^{l}$ satisfies $\|f\|^{2}_{H^{2}(\mathbb{D}^{2})}=\sum_{k,l}|a_{k,l}|^{2}<\infty$. Properties of Hardy spaces in the bidisk, and more generally the polydisk, are discussed in [34]. In particular, $H^{2}(\mathbb{D}^{2})$ can be identified via Cauchy integrals with the subspace of $L^{2}(\mathbb{T}^{2})$ whose elements have Fourier series with all nonnegative frequencies [34, Chapter 3]. On $H^{2}(\mathbb{D}^{2})$ we have a pair of shift operators
@@
If $\phi\colon\mathbb{D}^{2}\to\mathbb{C}$ is an inner function, that is, a bounded holomorphic function on the bidisk with unimodular non-tangential limits on $\mathbb{T}^{2}$, then the closed subspace $\phi H^{2}(\mathbb{D}^{2})$ is jointly invariant for $(T_{1},T_{2})$. In stark contrast to the one-variable situation, this does not describe all invariant subspaces (a subspace of $H^{2}(\mathbb{D}^{2})$ is of Beurling type when the shifts satisfy an additional doubly commuting condition [29]) and the operator theory in $H^{2}(\mathbb{D}^{2})$ is much more complicated. Function theory in the bidisk is similarly complicated: even rational inner functions $\phi=q/p$ need not extend to be continuous on $\mathbb{T}^{2}$ and can in fact have complicated singularities [11]. Thus, one expects that the structure of two-variable model spaces
@@
should be fairly complicated, see eg. [8] and references provided there.
@@
Nevertheless, there has been considerable recent progress on function theory in polydisks and associated operator theory. We do not give a comprehensive overview here, we merely focus on developments that are particularly relevant to the results in the present paper. Inspired by work of E. Doubtsov [20], Clark measures on $\mathbb{T}^{2}$ associated with rational inner functions have been studied in great detail [9, 4, 15], and detailed structure formulas for these measures have been obtained. See [14, 5, 24] for further extensions.
@@
In this paper, we investigate generalizations of Clark theory to two variables by studying commuting Clark unitaries $(U^{1}_{\alpha},U^{2}_{\alpha})$ obtained from compressions of the shifts $(T_{1},T_{2})$ to two-variable model spaces. [^1] Terminology varies across the literature, but in this paper, we take ‘Clark unitary (operator)’ to mean a perturbation of a compressed shift on a model space into a unitary operator. We will also encounter the unitary operators that embed model spaces in certain $L^{2}$ spaces; we will refer to such operators as ‘Clark embedding operators’. We caution that in two or more variables, these embedding operators are isometric but not always surjective! Our goals include giving a description of such unitaries and showing that, at least in the case of rational inner (RIFs) $\phi$, the Taylor joint spectrum of the unitaries (defined below) is equal to the support of the Clark measure $\sigma_{\alpha}$, a one-dimensional subset of the two-torus. Since several methods for constructing RIFs with prescribed properties are known, our results produce a wide range of examples of sets that can arise as Taylor spectra for natural operators acting on spaces of holomorphic functions on the bidisk.
@@
As is to be expected, there are several obstacles we need to overcome along the way. First is the realization that, in contrast to the classical one-dimensional case, we cannot expect finite rank perturbations of compressed shifts to unitaries. The perturbations we do obtain involve projections onto a model space which are not always easy to compute explicitly in closed form. The structure of RIF Clark measures is also fairly complicated: these measures act by integrating along unions of smooth curves in $\mathbb{T}^{2}$ and weighting by non-constant weights obtained from partial derivatives of, in general, globally discontinuous RIFs. Finally, we need to connect the properties of these objects with the structure of the chain complex defining the Taylor joint spectrum.
@@
We should emphasize that we focus on Clark theory in two variables in the setting of the bidisk; see for instance [25, 26, 2, 3] for the setting of the unit ball. There are several reasons for restricting our attention to the case of two variables. On the one hand, the function theory of rational inner functions is more complicated in three or more variables, one manifestation being that the level sets of a RIF in higher dimensions cannot, in general, be parameterized by analytic or even continuous functions [12]. Our understanding of Clark measures in higher-dimensional polydisks is less detailed (see however [15]), with examples of pathological behavior, such as having unbounded weight functions. On the other hand, the formalism of the Koszul complex [36] used to define the Taylor joint spectrum becomes more involved as the number of variables goes up, and well-known operator theoretic obstacles such as the failure of Andô’s inequality [31, 37] further support the investigations of pairs of commuting unitaries separately.
@@ 1.2. Overview
Our paper is structured as follows. In Section 2, we examine the simple example $\phi=z_{1}z_{2}$ and its associated Clark measures and Clark unitary operators. This example is simple enough to be handled from first principles, but the features we observe illustrate general results we obtain later on. Section 3 contains background material we use freely in the remainder of the paper. We review some salient results from Clark theory in one variable, before moving on to a discussion of model spaces and Clark measures in two variables with a special emphasis on the case of rational inner functions. Finally, we recall the definition of Taylor joint spectrum and list some of the features of this notion of multivariate spectrum. Next, in Section 4, under assumptions including the case of rational inner functions, we define pairs of Clark unitary operators $(U^{1}_{\alpha},U^{2}_{\alpha})$ starting with two-variable compressed shifts, after first computing the adjoint of Clark embedding operators. We show that these Clark operators on the model space $K_{\phi}$ are unitarily equivalent to multiplication by the coordinate functions on $L^{2}(\sigma_{\alpha})$. Under additional assumptions on the underlying inner function $\phi$, we obtain concrete formulas for Clark unitaries $(U^{1}_{\alpha},U^{2}_{\alpha})$. Section 5 identifies the Taylor joint spectrum of a pair of Clark unitaries with the corresponding $\alpha$-level set of the underlying inner function $\phi$ under some technical assumptions on the associated Clark measure $\sigma_{\alpha}$. This covers, in particular, the case where $\phi$ is rational inner. Several examples that illustrate results and obstructions are presented in the closing Section 6.
@@ 2. A Simple Example
$K_{\phi}=K_{z_{2}}\oplus z_{2}H^{2}_{2},$
where $K_{z_{2}}=H^{2}(\mathbb{D}^{2})\ominus z_{2}H^{2}(\mathbb{D}^{2})$ and $H^{2}_{2}$ denotes the one-variable Hardy space $H^{2}(\mathbb{D})$, embedded in $H^{2}(\mathbb{D}^{2})$ with independent variable $z_{2}$. We can readily describe the action of compressed shifts on $K_{z_{2}}$ and $z_{2}H^{2}_{2}$. We compute
$(S_{\phi}^{1}z_{1}^{k})(z)=(P_{K_{\phi}}T_{1}z_{1}^{k})(z)=(P_{K_{\phi}}z_{1}^{k+1})(z)=z_{1}^{k+1},\quad k=0,1,\ldots,$
@@
$S_{\phi}^{2}z_{1}^{k}=0,\quad k=0,1,\ldots,\quad\textrm{and}\quad S_{\phi}^{2}z_{2}^{l}=z_{2}^{l+1},\quad l=1,2,\ldots.$
This means that $S^{1}_{\phi}$ maps the subspace $K_{z_{2}}\subset K_{\phi}$ into itself and $z_{2}H^{2}_{2}$ onto $\{0\}$; an analogous statement holds for $S^{2}_{\phi}$. From this, it is clear that, in contrast to the classical one-variable setting, there are no finite-rank perturbations of $S^{1}_{\phi}$ and $S^{2}_{\phi}$ that result in unitary operators on $K_{\phi}$. For instance, note that $\mathrm{span}(\{1\}\cup\{z_{2}^{l}\colon l\geq 1\})\perp\mathrm{Ran}S^{1}_{\phi}$.
However, if we allow infinite-rank perturbations we can write down a concrete representation of unitaries associated with the compressed shifts. Namely, let $\alpha\in\mathbb{T}$ and consider the operators $R_{\alpha}^{k}\colon K_{\phi}\to K_{\phi}$, $k=1,2$ given by
$(R_{\alpha}^{k}f)(z)=\alpha\sum_{j=0}^{\infty}\left\langle f(w_{1},w_{2}),w_{k}^{j+1}\right\rangle z_{k}^{j},\quad k=1,2.$
Then, since $K_{z_{2}}\perp z_{2}H^{2}_{2}$, we have $R_{\alpha}^{1}(z_{1}^{k})=R_{\alpha}^{2}(z_{2}^{l})=0$ for all $k,l$ while
$R_{\alpha}^{1}(z_{2}^{l})=\alpha z_{2}^{l-1},\quad l=1,2,\ldots\qquad\textrm{and}\quad R_{\alpha}^{2}(z_{1}^{k})=\alpha z_{1}^{k-1},\quad k=0,1,\ldots.$
A more compact expression for $R_{\alpha}^{1}$, amenable to generalization, is
$(R_{\alpha}^{1}f)(z)=\alpha\left\langle f(w),T_{1}^{*}\phi(w)\frac{1}{1-\overline{z}_{2}w_{2}}\right\rangle_{H^{2}(\mathbb{D}^{2})},$
where $T_{1}^{*}$ denotes the adjoint of $T_{1}$; a similar expression is valid for $R_{\alpha}^{2}$.
Now consider the operators $U^{j}_{\alpha}\colon K_{\phi}\to K_{\phi}$, $j=1,2$, defined by
@@
Then by the above computations, $(U^{1}_{\alpha},U^{2}_{\alpha})$ form a pair of commuting unitaries on $K_{\phi}$ for each $\alpha\in\mathbb{T}$. Note that the splitting $K_{\phi}=K_{z_{2}}\oplus z_{2}H^{2}_{2}$ is not preserved. For instance, we have $U_{\alpha}^{1}(z_{2})=\alpha\in K_{z_{2}}$ while $z_{2}\in z_{2}H^{2}_{2}$; this is because $R_{\alpha}^{1}$ maps $z_{2}\in z_{2}H^{2}_{2}$ to $\alpha\in K_{z_{2}}$.

### v2

@@ Abstract.
We develop a Clark theory for commuting compressed shift operators on model spaces $K_{\phi}$ associated with inner functions $\phi$ on the bidisk, which exhibits both similarities and marked differences compared to the classical one-variable version. We first identify the adjoint of the embedding operator $J_{\alpha}\colon K_{\phi}\to L^{2}(\sigma_{\alpha})$ as a weighted Cauchy transform of the Clark measure $\sigma_{\alpha}$. Under natural assumptions, which generically include the case when $\phi$ is rational inner, we obtain commuting unitaries on $K_{\phi}$ that are (often infinite-dimensional) perturbations of the compressed shift operators $K_{\phi}$. We prove that these unitaries are unitarily equivalent to multiplication by the coordinate functions on $L^{2}(\sigma_{\alpha})$ and then establish a number of related properties and simplified results in special cases. Finally, we show that the Taylor joint spectrum of these Clark unitaries coincides with level sets of $\phi$ when $\phi$ is a rational inner function.
@@ 1.1. Introduction
is a fundamental example of an isometry. It is defined on holomorphic functions $f=\sum_{k=0}^{\infty}a_{k}z^{k}$ in the disk satisfying $\|f\|^{2}_{H^{2}(\mathbb{D})}=\sum_{k}|a_{k}|^{2}<\infty$ as $T\colon f\mapsto z\cdot f$. The shift operator has been studied extensively over the last century, and many important theorems regarding the Hardy shift have become classical results in operator theory (see, for instance, [28, Chapters 3 and 4]). An important example is a well-known theorem of Beurling which characterizes the invariant subspaces of the shift. Namely, $\mathcal{M}\subset H^{2}(\mathbb{D})$ is a non-trivial closed invariant subspace for $T$ if and only if $\mathcal{M}=\phi H^{2}(\mathbb{D})$ for some inner function $\phi\colon\mathbb{D}\to\mathbb{C}$. Recall that $\phi$ is inner if $\phi\in H^{\infty}(\mathbb{D})$, the Banach algebra of bounded holomorphic functions in the disk, and its non-tangential limits satisfy $|\phi^{*}(\zeta)|=1$ for almost every $\zeta$ on the unit circle
@@
$S_{\phi}=P_{{\phi}}T,$
where $P_{{\phi}}\colon H^{2}(\mathbb{D})\to K_{\phi}$ denotes orthogonal projection. The ‘model’ terminology arises from Szőkefalvi-Nagy-Foiaş theory, which realizes a wide class of Hilbert space contractions as compressed shifts on model spaces [28, Chapter 9]. This fact provides a powerful motivation for studying compressions of the shift acting on model spaces, and also for investigating how properties of the underlying inner function $\phi$ are reflected in its operator theory. An important contribution in this direction was made by D.N. Clark, who observed [23] that the rank-one perturbations of $S_{\phi}$ to a unitary operator form a one-parameter family $\{U_{\alpha}\}$ indexed by $\alpha\in\mathbb{T}$. He further identified the spectral measures of these unitaries as the positive singular Borel measures $\{\sigma_{\alpha}\}$ that satisfy
(1.1) $\frac{1-|\phi(z)|^{2}}{|\alpha-\phi(z)|^{2}}=\int_{\mathbb{T}}P_{z}(\zeta)d\sigma_{\alpha}(\zeta)$
where $P_{z}(\zeta)=(1-|z|^{2})/|\zeta-z|^{2}$ denotes the Poisson kernel of the disk. These measures are often referred to as Clark measures , and their properties are intimately connected, through a unitary map $J_{\alpha}\colon K_{\phi}\to L^{2}(\sigma_{\alpha})$, with both the function theory of $\phi$ and the spectral theory of $U_{\alpha}$. For instance, $\sigma_{\alpha}$ is the spectral measure of $U_{\alpha}$, and a point in $\mathrm{supp}\,\sigma_{\alpha}\subset\mathbb{T}$ is an eigenvalue of $U_{\alpha}$ precisely if $\phi$ has a finite angular derivative at the corresponding point of the unit circle. We refer the reader to [22, 28] for accessible introductions to classical Clark theory, and to [38, 40, 27, 34] for a more varied accounting of recent progress (eg. including the case of non-inner $\phi$).
@@
and its distinguished boundary $\mathbb{T}^{2}=\{(\zeta_{1},\zeta_{2})\in\mathbb{C}^{2}\colon|\zeta_{j}|=1,\,j=1,2\}$. We say that a holomorphic function $f\colon\mathbb{D}^{2}\to\mathbb{C}$ belongs to $H^{2}(\mathbb{D}^{2})$ if $f=\sum_{k=0}^{\infty}\sum_{l=0}^{\infty}a_{k,l}z_{1}^{k}z_{2}^{l}$ satisfies $\|f\|^{2}_{H^{2}(\mathbb{D}^{2})}=\sum_{k,l}|a_{k,l}|^{2}<\infty$. Properties of Hardy spaces in the bidisk, and more generally the polydisk, are discussed in [41]. In particular, $H^{2}(\mathbb{D}^{2})$ can be identified via Cauchy integrals with the subspace of $L^{2}(\mathbb{T}^{2})$ whose elements have Fourier series with all nonnegative frequencies [41, Chapter 3]. On $H^{2}(\mathbb{D}^{2})$ we have a pair of shift operators
@@
If $\phi\colon\mathbb{D}^{2}\to\mathbb{C}$ is an inner function, that is, a bounded holomorphic function on the bidisk with unimodular non-tangential limits on $\mathbb{T}^{2}$, then the closed subspace $\phi H^{2}(\mathbb{D}^{2})$ is jointly invariant for $(T_{1},T_{2})$. In stark contrast to the one-variable situation, this does not describe all invariant subspaces (a subspace of $H^{2}(\mathbb{D}^{2})$ is of Beurling type when the shifts satisfy an additional doubly commuting condition [35]) and the operator theory in $H^{2}(\mathbb{D}^{2})$ is much more complicated. Function theory in the bidisk is similarly complicated: even rational inner functions $\phi=q/p$ need not extend to be continuous on $\mathbb{T}^{2}$ and can in fact have intricate singularities [17]. Thus, one expects that the structure of two-variable model spaces
@@
(where $P_{{\phi}}\colon H^{2}(\mathbb{D}^{2})\to K_{\phi}$ denotes orthogonal projection) should be fairly complicated, see eg. [13, 9] and references provided there.
@@
Nevertheless, there has been considerable recent progress on function theory in polydisks and associated operator theory. We do not give a comprehensive overview here, we merely focus on developments that are particularly relevant to the results in the present paper. Model spaces (also called Beurling quotient modules) have recently been characterized in terms of certain operator-theoretic conditions in [9, Theorem 1.1]. Inspired by work of E. Doubtsov [26], Clark measures on $\mathbb{T}^{2}$ associated with rational inner functions have been studied in great detail [14, 6, 21], and detailed structure formulas for these measures have been obtained. See [20, 7, 30] for further extensions.
@@
In this paper, we investigate generalizations of Clark theory to two variables by studying commuting Clark unitaries $(U^{1}_{\alpha},U^{2}_{\alpha})$ obtained from compressions of the shifts $(T_{1},T_{2})$ to two-variable model spaces. [^1] Terminology varies across the literature, but in this paper, we take ‘Clark unitary (operator)’ to mean a perturbation of a compressed shift on a model space into a unitary operator. We will also encounter the unitary operators that embed model spaces in certain $L^{2}$ spaces; we will refer to such operators as ‘Clark embedding operators’. We caution that in two or more variables, these embedding operators are isometric but not always surjective! Our goals include giving a description of such unitaries, investigating their properties, and showing that, at least in the case of rational inner function (RIF) $\phi$, the Taylor joint spectrum of the unitaries (defined below) is equal to the support of the Clark measure $\sigma_{\alpha}$, a one-dimensional subset of the two-torus. Since several methods for constructing RIFs with prescribed properties are known, our results produce a wide range of examples of sets that can arise as Taylor spectra for natural operators acting on spaces of holomorphic functions on the bidisk.
@@
As is to be expected, there are several obstacles we need to overcome along the way. First is the realization that, in contrast to the classical one-dimensional case, we cannot expect finite rank perturbations to be sufficient to turn compressed shifts to unitaries. The perturbations we do obtain involve projections onto the model spaces, which are not always easy to compute explicitly in closed form. The structure of RIF Clark measures is also fairly complicated: these measures act by integrating along unions of smooth curves in $\mathbb{T}^{2}$ and weighting by non-constant weights obtained from partial derivatives of, in general, globally discontinuous RIFs. Finally, we need to connect the properties of these objects with the structure of the chain complex defining the Taylor joint spectrum.
@@
We should emphasize that we focus on Clark theory in two variables in the setting of the bidisk; see for instance [31, 32, 3, 4] for the setting of the unit ball. There are several reasons for restricting our attention to the case of two variables. On the one hand, the function theory of rational inner functions is more complicated in three or more variables, one manifestation being that the level sets of a RIF in higher dimensions cannot, in general, be parameterized by analytic or even continuous functions [18]. Our understanding of Clark measures in higher-dimensional polydisks is less detailed (see however [21]), with examples of pathological behavior, such as the measures having unbounded weight functions. On the other hand, the formalism of the Koszul complex [42] used to define the Taylor joint spectrum becomes more involved as the number of variables increases and well-known operator theoretic obstacles such as the failure of Andô’s inequality [37, 43] further support the investigations of pairs of commuting unitaries separately.
@@ 1.2. Overview
Our paper is structured as follows. In Section 2, we study the simple example $\phi=z_{1}z_{2}$ and its associated Clark measures and Clark unitary operators. This example is simple enough to be handled from first principles, but the features we observe illustrate the general results that we obtain later on.
In Section 3, we present background material that we use freely in the remainder of the paper. We first review some salient results from Clark theory in one variable and then move to a discussion about model spaces (including related Agler decompositions) and Clark measures in two variables, with a special emphasis on the case of rational inner functions. Finally, we recall the definition of Taylor joint spectrum and list some of the features of this notion of multivariate spectrum.
In Section 4, we investigate the existence and properties of Clark operators on the bidisk. We start by considering the Clark embedding operators from the model space $K_{\phi}$ into the Clark space $L^{2}(\sigma_{\alpha})$ and identifying their adjoints. Then in Theorem 4.3, under further assumptions that the embedding operator is unitary and related functions $\vartheta^{1}_{\alpha},\vartheta^{2}_{\alpha}$ are bounded, we prove that the operators $(U_{\alpha}^{1},U_{\alpha}^{2})$ defined on the associated model space by
give perturbations of the compressed shifts $(S_{\phi}^{1},S_{\phi}^{2})$ to commuting unitary operators. Here, $M_{\overline{\vartheta^{j}_{\alpha}}}$ is multiplication by $\overline{\vartheta^{j}_{\alpha}}$ on $L^{2}$ and $P_{H^{2}_{j}}$ is the projection from $L^{2}$ onto a one-variable Hardy space. As in the one-variable situation, we prove that these Clark operators $(U_{\alpha}^{1},U_{\alpha}^{2})$ are unitarily equivalent to multiplication by the coordinate functions on $L^{2}(\sigma_{\alpha})$.
The remainder of Section 4 investigates special cases and properties of $(U_{\alpha}^{1},U_{\alpha}^{2})$. For example, Corollary 4.5 establishes very simple formulas for these operators in the setting where $\phi$ has a factor of $z_{1}z_{2}.$ Meanwhile, Corollary 4.6 shows that the assumptions from Theorem 4.3 generically hold when $\phi$ is a rational inner function of two variables. Proposition 4.7 shows that the projection $P_{\phi}$ is usually needed in the $U_{j}^{\alpha}$ formulas and Theorem 4.9 shows that, while these perturbations of compresssed shifts to unitaries are often infinite-dimensional, they are still “small” in a way that naturally aligns with the known structure of the two-variable model spaces.
Then, Section 5 studies the Taylor joint spectrum of pairs of Clark unitaries. Under the assumption that $\phi$ is a two-variable rational inner function without a monomial factor, Theorem 5.4 shows that the Taylor joint spectrum of $(U_{\alpha}^{1},U_{\alpha}^{2})$ is the $\alpha$ level set of $\phi$ on $\mathbb{T}^{2}$. As part of this analysis, we exploit the precise structure of Clark measures of rational inner functions.
In Section 6, we illustrate our results by applying them to both products of one-variable finite Blaschke products and a singular rational inner function. In both settings, our computations illustrate obstructions to finding tractable ways of explicitly computing $(U_{\alpha}^{1},U_{\alpha}^{2})$ when $\phi$ is not divisible by $z_{1}z_{2}.$
@@ 2. A Simple Example
$K_{\phi}=H^{2}_{1}\oplus z_{2}H^{2}_{2}=H^{2}_{2}\oplus z_{1}H^{2}_{1},$
where for $j=1,2$, $H^{2}_{j}$ denotes the one-variable Hardy space $H^{2}(\mathbb{D})$, embedded in $H^{2}(\mathbb{D}^{2})$ with independent variable $z_{j}$. We can readily describe the action of compressed shifts on powers of $z_{j}$:
$(S_{\phi}^{1}z_{1}^{k})(z)=(P_{{\phi}}T_{1}z_{1}^{k})(z)=(P_{{\phi}}z_{1}^{k+1})(z)=z_{1}^{k+1},\quad k=0,1,\ldots,$
@@
$S_{\phi}^{2}z_{1}^{k}=0,\quad k=1,2\ldots,\quad\textrm{and}\quad S_{\phi}^{2}z_{2}^{l}=z_{2}^{l+1},\quad l=0,1,2,\ldots.$
This means that $S^{1}_{\phi}$ maps the subspace $H^{2}_{1}\subseteq K_{\phi}$ into itself and $z_{2}H^{2}_{2}$ onto $\{0\}$; an analogous statement holds for $S^{2}_{\phi}$. From this, it is clear that, in contrast to the classical one-variable setting, there are no finite-rank perturbations of $S^{1}_{\phi}$ and $S^{2}_{\phi}$ that result in unitary operators on $K_{\phi}$. For instance, note that $H^{2}_{2}\subseteq K_{\phi}$ and $H^{2}_{2}\perp\mathrm{Ran}S^{1}_{\phi}$.
However, if we allow infinite-rank perturbations we can write down a concrete representation of unitaries associated with the compressed shifts. Namely, let $\alpha\in\mathbb{T}$ and consider the operators $R_{\alpha}^{k}\colon K_{\phi}\to K_{\phi}$, $k=1,2$ defined on powers of $z_{j}$ by $R_{\alpha}^{1}(z_{1}^{k})=R_{\alpha}^{2}(z_{2}^{l})=0$ for all $k,l\geq 0$ and
$R_{\alpha}^{1}(z_{2}^{l})=\alpha z_{2}^{l-1},\quad l=1,2,\ldots\qquad\textrm{and}\quad R_{\alpha}^{2}(z_{1}^{k})=\alpha z_{1}^{k-1},\quad k=1,2,\ldots.$
A more compact expression for the $R_{\alpha}^{j}$ operators, amenable to generalization, is
$R_{\alpha}^{1}=\alpha P_{H_{2}^{2}}M_{\bar{\phi}/\bar{z_{1}}}\ \text{ and }\ \ R_{\alpha}^{2}=\alpha P_{H^{2}_{1}}M_{\bar{\phi}/\bar{z_{2}}}$
where $P_{H^{2}_{j}}$ is the orthogonal projection from $L^{2}$ onto $H^{2}_{j}$ and $M_{\bar{\phi}/\bar{z_{j}}}$ is multiplication by the bounded function $\bar{\phi}/\bar{z_{j}}$ on $L^{2}$. Now consider the operators $U^{j}_{\alpha}\colon K_{\phi}\to K_{\phi}$, $j=1,2$, defined by
@@
Then if $f(z)=f_{1}(z_{1})+z_{2}f_{2}(z_{2})\in K_{\phi}$ for $f_{1}\in H^{2}_{1},f_{2},\in H^{2}_{2},$ then
$(U^{1}_{\alpha}f)(z)=z_{1}f_{1}(z_{1})+\alpha f_{2}(z_{2}),$
and a similar formula holds for $U^{2}_{\alpha}.$ From this formula, one can see that $U^{1}_{\alpha}$ is unitary and with direct computations, one can easily check that the pair of unitaries $(U^{1}_{\alpha},U^{2}_{\alpha})$ commute. Note that the splitting $K_{\phi}=H^{2}_{1}\oplus z_{2}H^{2}_{2}$ is not preserved. For instance, we have $U_{\alpha}^{1}(z_{2})=\alpha\in H_{1}^{2}$ while $z_{2}\in z_{2}H^{2}_{2}$; this is because $R_{\alpha}^{1}$ maps $z_{2}\in z_{2}H^{2}_{2}$ to $\alpha\in H^{2}_{1}$. These formulas for $R_{\alpha}^{j}$ and $U_{\alpha}^{j}$ align with the more general perturbations of compressed shifts to unitaries that will appear later in Theorem 4.3 and Corollary 4.5.